OpenStudy (44691):

Solve 2x2 = 50. {±25} {±12.5} {±5} {±6}

rishavraj (rishavraj):

@44691 u know this one too ???? just asking

OpenStudy (44691):

yes lol. Do you?

rishavraj (rishavraj):

yeah obviously

OpenStudy (44691):

ok. ill explain so you know that I know.

rishavraj (rishavraj):

lol no need :)

OpenStudy (44691):

Step 1: Divide both sides by 2 \[\frac{ 2x^2 }{ 2 } = \frac{ 50 }{ 2 }\] Which equals: \[x^2 = 25\], Step 2. Take the Square Root of 25: \[x = \pm \sqrt{25}\] Which equals: \[x = 5 ~or~ x = -5\], so the answer would be \[Answer: \pm 5\]

OpenStudy (44691):

@rishavraj am I right?

rishavraj (rishavraj):

yup :)

OpenStudy (44691):

ok:) I'm normally right. I'm pretty good at math lol. :)